INAF queue observing, 2018 June 28-29

Observer: A. Rossi
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson (on call)
Instrument: LUCI1 and LUCI2


At the sunset the sky is half covered with clouds
instruments: MODS2 red grating is not working well
Night duration (12\xB0 el twilight) 3:35- 11:10 UT = 7 h 35m

Observing time: h m (%)

Weather time loss: h m (%)

technical time loss h (%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
CALIB DARKS LUCI1+LUCI2 H spectra     For Rossi 49
CALIB DARKS LUCI2 HK,zJ spectra     For Rossi 47
CALIB DARKS LUCI2 imaging     For Rossi 47
CALIB DARKS LUCI2 HK,zJ spectra     For Antoniucci V371Ser and V2492Cyg +telluric
STD HIP 102074   zJ HK   luci1.20180629.0024-0032, luci2.20180629.0042-0048 For Antoniucci V2492Cyg
Antoniucci 8 V2492Cyg LUCI1+LUCI2 zJ HK 1920/1920s luci1.20180629.0037-0044, luci2.20180629.0054-0061 long slit completed
STD   LUCI1+LUCI2     luci1.20180629.0048-0051, luci2.20180629.0065-68 telluric Polletta Phz1119
Polletta Phz1119 LUCI1+LUCI2 HKspec 4800/14400s luci1.20180629.0056-0075, luci2.20180629.0073-0092  
Rossi47 GRB 071021 LUCI2 HKonly 0.5/2h luci2.20180629.0101-0108 to be checked seeing jumping +cirrus
STD HIP111538 LUCI2 HK   luci2.20180629.0112-0115 for ROssi47 GRB071021
CALIB arc flats LUCI1+LUCI2 zJ HKspec     for Antoniucci V2492Cyg and V371Ser
CALIB DOME FLATS LUCI2 N30, H Ks     for Giannetti AO

  • luci1.20180629.0001-0005, luci2.20180629.0001-0005 DIT=150, NDIT=1 MER NORM for Rossi49
  • luci1.20180629.0006-0010, luci2.20180629.0006-0010 DIT=5, NDIT=6 LIR INT for Rossi49
  • luci2.20180629.0011-0015 DIT 5 NDIT 6 Rossi 47 telluric
  • luci2.20180629.0016-0020 DIT 2.5 NDIT 25 Rossi 47 imaging
  • luci2.20180629.0021-0025 DIT 210 NDIT 1 MER NORM Rossi 47 spectroscopy
  • luci1.20180629.0011-0015, luci2.20180629.0027-0031 DIT 240 NDIT 1 MER NORM Antoniucci V371Ser and V2492Cyg
  • luci1.20180629.0016-0020, luci2.20180629.0032-0036 DIT 2.75 NDIT 1 MER NORM Antoniucci V371Ser and V2492Cyg telluric

04:45 we open but wait for clouds to ,move, we cannot collimate on anything frown, sad smile


06:00 slew to telluric for Antoniucci V2492 cyg, HIP 102074
  • cirrus, seeing 1.2-1.3
  • acq luci1.20180629.0021-0023, luci2.20180629.0037-0039
  • 10 min loss camera on luci1 in unknown status error
  • 06:15 seeing jumpiong between 1" and 1.4"
  • sci luci1.20180629.0024-0032, luci2.20180629.0042-0048

06:29 slew to Antoniucci V2492 cyg zJ HK spec
  • cirrus, seeing 1.2-1.3
  • 5 min loss camera on luci1 in unknown status error
  • acq luci1.20180629.0033-0036, luci2.20180629.0049-0053
  • sci luci1.20180629.0037-0044, luci2.20180629.0054-0061
  • DIMM 1.3" at 07:20

07:21 slew to telluric for Polletta Phz 1119 HK spec
  • acq luci1.20180629.0045-0047, luci2.20180629.0062-0063
  • sci luci1.20180629.0048-0051, luci2.20180629.0064-68
  • 5minloss because Andrea does not remember if there was a problem with binocular MOS telluric acquisition. Indeed there is, see bottom of the page

:07:52 Slew to Polletta Phz1119 HK spec only
  • DIMM 1.3" at 07:55, some cirrus sometimes
  • acq luci1.20180629.0052-0055, luci2.20180629.0069-0072
  • sci luci1.20180629.0056-0075, luci2.20180629.0073-0092
  • DIMM 1.5" at 08:15
  • DIMM 1.2" at 08:35

09:05 Seeing stable at 1.1" try ESM Rossi
  • authorize LUCI2 only

09:20 AOworking, start script 071021 spettroscopy HK only
  • 30min loss MOS error start again at 09:55
  • clouds we stop
  • 10:20 try again
  • acq luci2.20180629.0097-0100
  • sci luci2.20180629.0101-0108 first 2 exposure do not useful, seeing at 1.2"
  • seeing jump to 1.5"

10:05 Slew to telluric HIP111538 HK only
  • acq luci2.20180629.0109-0111
  • sci luci2.20180629.0112-0115

  • luci1.20180629.076-0091, luci2.20180629.0116-0129 zJ arcflats 0.75" 1.175um for Antoniucci V2492Cyg and V371Ser
  • luci1.20180629.092-0105, luci2.20180629.0130-0245 HK arcflats 0.75" 1.93um for Antoniucci V2492Cyg and V371Ser
  • luci2.20180629.0173-0182 DOME FLATS N30 H-band Giannetti
  • luci2.20180629.0183-0192 DOME FLATS N30 Ks-band Giannetti

NOTE: for telluric MOS. When yo calculated the offsets, do not send them!!
  • put them in the telsvc GUI (the one where you can write the offset)
  • click on left offset and right offset. They have to be green
  • click on sync offset
  • click commit and wait the 2 happy sounds!
  • click absorb
  • now go to LUCI GUI and jump to the next step, so you jump the step when offsets are given
  • click go and check if you get spectra in the slit on the left


CALIB DARKS LUCI1+LUCI2 DIT=120s Mer Norm     For PollettaPhz1119
| CALIB | N30 SKYFLATS | LUCI2 | H, Ks ima | | | For Giannetti | but there are domeflats taken on 20180629

CALIB MOS arcs+flats LUCI1+LUCI2 HKspec 1.93     For PollettaPhz1119
-- %USERSIG{AndreaRossi - 2018-06-28}%


Topic revision: r7 - 30 Jun 2018, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback